OpenStudy (brydenwright):

Use the angle addition formula to find the expression for cos (a+b). What happens if you change it to cos(a+a)

OpenStudy (legomyego180):

Heres a hint: \[Cos(a+a)=Cos(2a)\]

TheSmartOne (thesmartone):

Hint: cos(A+B) = cosAcosB − sinAsinB